OpenStudy (anonymous):

x 2 + y 2 + 6x − 8y + 8 = 0 represents a circle. Find its centre and show that its radius is √ 17.

OpenStudy (anonymous):

center is (-3,4) and radius is \[\sqrt{17}\]

OpenStudy (anonymous):

many thanks. please could you show me how you got this if possible

OpenStudy (anonymous):

circle: (x-a)^2 +(y-b)^2=R^2 <=> x^2+y^2-2ax-2yb+a^2+b^2-R^2=0

OpenStudy (anonymous):

with center (a,b) radius R

OpenStudy (anonymous):

in case: -2a=6 ->a=-3 -2b=-8 ->b=4 a^2+b^2-R^2=8 -> R=sqrt(17)

OpenStudy (anonymous):

thank you that makes some sense. thank you for your help